OpenStudy (anonymous):

How many possibilities are there if you are creating a 4-digit password using the digits 0 through 9? You may repeat numbers. A.) 40 B.) 6,561 C.) 5,040 D.) 10,000

OpenStudy (anonymous):



OpenStudy (reemii):

(number of posibilities for 1st digit) \(\times\)(number of posibilities for 2nd digit) \(\times\)(number of posibilities for 3rd digit) \(\times\)(number of posibilities for 4th digit) First the 1st digit there are 10 choices, what about the others?

OpenStudy (anonymous):